OpenStudy (anonymous):

Simplify expression: (5+x^-1)/(25-x^-2)

OpenStudy (anonymous):

x/(5x-1)

OpenStudy (anonymous):

you can use pure algebra with exponential n http://www.wolframalpha.com/otation, but if you really want to see what i going on, recall that \(b^{-n}=\frac{1}{b^n}\) and multiply \[(5+\frac{1}{x})(25-\frac{1}{x^2})\]

OpenStudy (anonymous):

How exactly do you solve it? Could you explain how you got the answer?

OpenStudy (anonymous):

you need to do four multiplications, just as you would if you had \((x+3)(x+4)\)

OpenStudy (anonymous):

use x^-1=1/x and then you have (5+1/x)/(25-1/x^2)

OpenStudy (anonymous):

(5x+1)/ x/ (25x^2-1) /(x^2) this is equal to (5x+1)x^2/x(5x-1)(5x+1) wich is equal to x/(5x-1)

OpenStudy (anonymous):